# biblio.bib

### v1

@@ A.1 Generative model sampling modes
Here, we show how to define the generative model $p_{\theta}(\mathbf{h}_{0:1},\mathbf{w}_{0:1}|\mathbf{c})$ of Eq. 4, using the appropriate reverse transitions stemming from Eq. 5 and Eq. 6, as explained in the main text. For the sake of clarity, we drop the condition on $\mathbf{c}$.
@@
    Sequential modes: Equivalently to Eq. 14, we can sample first the discrete tokens $\mathbf{h}_{0}$ independently, and sequentially generate the corresponding parameters $\mathbf{w}_{0}$ by
@@
    Single mode: Building upon Eq. 15, we can also assume that we are given a gate Ansatz, which means a given, fixed $\mathbf{h}_{0}$. If we want to generate the corresponding parameters, we can then just use the second part of previous equation, namely,
@@
    Independent modes: Finally, the use of classifier-free guidance (CFG) [ho2022classifier] requires the definition of unconditional marginals, i.e. $p(\mathbf{h}_{0})$ and $p(\mathbf{w}_{0})$, which here we can sampled using
@@
As (ii) requires double the number of model evaluations, we only use (i) for joint sampling in Section 4.
@@ A.2 Multimodal top-down posterior
We show here that the top-down posterior from Eq. 5 splits to Eq. 6. Considering $0\leq s<t\leq 1$ and $0\leq\tilde{s}<\tilde{t}\leq 1$, we analyze the multimodal posterior by splitting the joint distribution
@@
In the previous equation, the second equality was done by means of the forward process Eq. 3, that allows us to derive
@@
A similar derivation can be done for the term $q({\color[rgb]{0,1,1}\mathbf{h}_{t}}|\mathbf{h}_{s},\mathbf{w}_{\tilde{t}},\mathbf{x})$ in the denominator in Eq. 20. We are using similar variants of Section A.2 below, i.e. for $q({\color[rgb]{0,.5,.5}\mathbf{h}_{s}}|{\color[rgb]{1,0,1}\mathbf{w}_{\tilde{s}}},\mathbf{w}_{\tilde{t}},\mathbf{x})$ and $q({\color[rgb]{0,.5,.5}\mathbf{h}_{s}}|\mathbf{w}_{\tilde{t}},\mathbf{x})$ in Eq. 22, and for $q({\color[rgb]{0.76,0.01,0}\mathbf{w}_{\tilde{t}}}|\mathbf{h}_{t},{\color[rgb]{0,.5,.5}\mathbf{h}_{s}},\mathbf{x})$ and $q({\color[rgb]{0.76,0.01,0}\mathbf{w}_{\tilde{t}}}|\mathbf{h}_{t},\mathbf{x})$ in Eq. 23 . Further, we simplify Eq. 20 by
@@
which is the usual top-down posterior for $\mathbf{w}_{\tilde{s}}$ for standard Gaussian DMs, as used for instance in DDPM [ho2020denoising]. In the same direction as done above, for the top-down posterior of $\mathbf{h}_{s}$ (the first part of Eq. 19) we derive
@@
The Eq. 24 proves the splitting of the reverse transitions of Eq. 6. Specifically, we set for the model
@@ A.3 Multimodal ELBO
which reduces to the loss of the main paper (see Eq. 7) when accounting for: 1) all sampling modes from Section A.1; 2) the use of the multimodal top-down posterior of Section A.2.
@@ A.4 Velocity loss weight
As described in Section 3.1, we use the velocity target [salimansprogressive]
@@
ensuring that the model outputs have unit variance. In contrast, the rectified flow-based velocity [liu2023flow], defined as $\mathbf{v}_{\text{flow}}\coloneq{\bf\it\epsilon}-\mathbf{x}$, does not guarantee this property. For a Gaussian diffusion process (see Eq. 2), we sample a latent $\mathbf{z}_{t}$ at time $t$ using
@@
where ${\bf\it\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Next, we combine Eq. 28 and Eq. 29 to
@@
Next, we derive the conversion factor $\omega_{x}(t)=(\mathrm{SNR}(t)+1)\,\omega_{v}(t)$, mentioned in Section 3.1, that arises when one translates the loss from a denoising prediction $\hat{\mathbf{x}}_{\theta}$ to a velocity predictor $\hat{\mathbf{v}}_{\theta}$. For this, we insert Eq. 30 into the MSE
@@
:-$\coloneq\omega_{v}(t)\norm{\left(\mathbf{v}_{t}-\hat{\mathbf{v}}_{\theta}\right)}_{2}^{2},:$ (31)
@@
which we use for the loss weights in Eq. 7, as introduced in the main text.
@@ A.5 Signal-to-noise ratio for continuous parameters
where $\bar{\beta}_{t}^{w}=1-\bar{\alpha}_{t}^{w}$. Moreover, we consider an embedding of the parameter $\lambda\in[-1,1]$ via (see Section 3.2 and Section B.2)
@@
As explained in Section 3.1, we use a sigmoid log-SNR weighting from Ref. [hoogeboom2024simpler]
@@
Notably, due to the encoding of the parameter $\lambda$ (see Eq. 34 and Section B.2), we have to alter the normal SNR in Eq. 35, initially given by $\mathrm{SNR}_{w}(t)\coloneq\bar{\alpha}_{t}^{w}/(1-\bar{\alpha}_{t}^{w})$. This redefinition is required because the signal is now encoded into two correlated dimensions, i.e. $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, and is therefore more robust to the noise of Eq. 33. The new $\mathrm{SNR}_{\lambda}$ for the continuous parameters $\lambda$, with its noisy reconstruction $\hat{\lambda}_{t}$ (see Section B.2), is defined as
@@
Here, we assume that the samples of $\lambda$ from the prior $p(\lambda)$ are bounded and normalized. Since $\lambda$ is periodic in $[-1,1]$ the absolute value is not relevant and we treat $\operatorname{\mathbb{E}}\left[\lambda^{2}\right]$ as a constant, as it does not depend on the diffusion time $t$. Indeed, we can view it as a constant shift for the sigmoid weighting in Eq. 35.
Given the previous definition, we can get an upper bound of Eq. 36 by using the Cramér–Rao bound [cramer1946mathematical] for an estimator of the parameter $\lambda$,
@@
where $\mathcal{I}(\cdot)$ is the Fisher information. We do so by first, calculating the score function $s(\lambda,\mathbf{w}_{t})$ of Eq. 33
@@
with evaluating the derivative of Eq. 34
@@
$=\frac{\bar{\alpha}_{t}^{w}}{\bar{\beta}_{t}^{w}}\pi^{2}\left(\sin(\lambda\pi)^{2}\norm{\mathbf{v}_{1}}^{2}+\cos(\lambda\pi)^{2}\norm{\mathbf{v}_{2}}^{2}-2\sin(\lambda\pi)\cos(\lambda\pi)\innerproduct{\mathbf{v}_{1}}{\mathbf{v}_{2}}\right)$ (48)
@@
where we used our orthogonal unit-variance normalized basis for $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ (see Section B.2). Given the previous, we define
@@
Therefore, for velocity prediction (see Section A.4), we set the weighting to
$\omega_{w}(t)$ :-$\coloneq(1-\bar{\alpha}_{t}^{w})\cdot\mathrm{sigmoid}[\log(\mathrm{SNR}_{\lambda}(t))]:$
@@ A.6 Learned noise schedule for discrete tokens
The problem of too-weak-noise schedulers for diffusing discrete token embeddings is visualized in Fig. 2a, e.g. for the cosine schedule [nichol2021improved], typically used for images. As explained in Section 3.2, we aim to learn a noise schedule with a desired average Hamming distance profile, such that the time of trivial denoising is minimized. Note that the Hamming distance for a single token can only be either 0 (not flipped) or 1 (flipped). Instead, when averaging w.r.t. to a sampling process, we can use the average Hamming distance, which is effectively the probability of a token flipping into any other token.
@@
$p_{\text{flip}}(t)$ $=1-\operatorname{\mathbb{E}}_{\mathbf{h}_{t}^{(i)}\sim q(\mathbf{h}_{t}^{(i)}|\mathbf{h}_{0}^{(i)})}\left[\mathrm{softmax}_{j}\left(\frac{1}{\tau}\innerproduct{\mathbf{h}_{0}^{(j)}}{\mathbf{h}_{t}^{(i)}}\right)\right]_{i}$
$=1-\operatorname{\mathbb{E}}_{\mathbf{h}_{t}^{(i)}\sim q(\mathbf{h}_{t}^{(i)}|\mathbf{h}_{0}^{(i)})}\left[\frac{\exp(\frac{1}{\tau}\innerproduct{\mathbf{h}_{0}^{(i)}}{\mathbf{h}_{t}^{(i)}})}{\sum_{j}\frac{1}{\tau}\innerproduct{\mathbf{h}_{0}^{(j)}}{\mathbf{h}_{t}^{(i)}}}\right],$ (53)
where $\tau>0$ is a temperature, which we generally set to $\tau=1/\sqrt{d_{h}}$ for a token embedding dimension of $d_{h}$. Analogous to the average Hamming distance, this probability $p_{\text{flip}}(t)$ is upper bounded by $1-1/N$, i.e., when $\mathbf{h}_{t}^{(i)}$ is sampled from the uniform distribution (see $t=1$ at Fig. 2a).
@@
In order to match $p_{\text{flip}}(t)$ to a desired Hamming distance target $f_{\text{target}}(t)$, we optimize the noise schedule $\bar{\alpha}_{t}^{h}$ appearing in Section A.6 by minimizing
@@
In practice, we found it easier to optimize $\bar{\alpha}_{t}^{h}$ indirectly by parameterizing it with $\alpha_{t}^{h}$, using the cumulative product relation $\bar{\alpha}_{t(i)}^{h}=\prod_{k=0}^{i}\alpha_{t(k)}^{h}$ (see Eq. 12 and Eq. 13) and treating $\alpha_{t(k)}^{h}$ as learnable parameters. Before optimization, we initialize the schedule values $\alpha_{t(k)}^{h}$ with the cosine schedule [nichol2021improved].
@@
In Fig. 2a we show the average Hamming distance of different optimized schedules according to Eq. 54. The curves plotted correspond to the following noise schedules:
@@
In Section 4, we use for all experiments a learned discrete noise schedule for the linear target (see Fig. 2a and c).
@@
As shown in Section 3.3, we use the duality between uniform discrete-state diffusion and continuous-state Gaussian diffusion [sahoo2025diffusionduality] to write the probability of finding a decoded, one-hot encoded token $\mathbf{k}_{t}\in\mathbb{R}^{N}$ for a discrete schedule $a_{t}$ as
@@
Just as the continuous SNR (see Eq. 2), we have the same boundary behavior of no noise at small diffusion times ($\lim_{t\rightarrow 0}\mathrm{SNR}_{\text{discrete}}(t)\rightarrow\infty$) to full noise at time $t=1$ ($\mathrm{SNR}_{\text{discrete}}(t=1)=0$). Here, full noise means a uniform distribution across the classes.
Similar to the continuous mode, we use the sigmoid log-SNR weighting of Eq. 35 (see Section 3.1). In order to account for the discrete nature of the embeddings, which embody high correlations across the embedding dimensions, we adjust the normal $\mathrm{SNR}_{h}(t)$ (Section 3.1) to the discrete one $\mathrm{SNR}_{\text{discrete}}(t)$ of Eq. 57.
To determine the appropriate weight $\omega_{h}(t)$ in Eq. 7, we relate the Gaussian flip probabilities $p_{\text{flip}}(t)$ (Section A.6) to the discrete schedule $a_{t}$ of Eq. 56. As all tokens are equivalently encoded, it suffices to look at only one token $i$. Hence, following Eq. 56, we write the probability of finding the token not to be flipped as
@@
resulting in the claimed relation of Section 3.3
@@
Therefore, assuming a well optimized learned schedule with Hamming target $f_{\text{target}}(t)$ (see Eq. 54), we can rewrite the discrete SNR as
@@
$\approx\frac{f_{\text{target}}(1)-f_{\text{target}}(t)}{f_{\text{target}}(t)},$ (60)
Finally, for the velocity prediction (see Section A.4), we set the weighting to
$\omega_{h}(t)$ :-$\coloneq(1-\bar{\alpha}_{t}^{h})\cdot\mathrm{sigmoid}[\log(\mathrm{SNR}_{\text{discrete}}(t))]:$

### v2

@@ A.1 Generative model sampling modes
Here, we show how to define the generative model $p_{\theta}(\mathbf{h}_{0:1},\mathbf{w}_{0:1}|\mathbf{c})$ of Eq.˜11, using the appropriate reverse transitions stemming from Eq.˜12 and Eq.˜13, as explained in the main text. For the sake of clarity, we drop the condition on $\mathbf{c}$.
@@
    Sequential modes: Equivalently to Eq.˜14, we can sample first the discrete tokens $\mathbf{h}_{0}$ independently, and sequentially generate the corresponding parameters $\mathbf{w}_{0}$ by
@@
    Single mode: Building upon Eq.˜15, we can also assume that we are given a gate Ansatz, which means a given, fixed $\mathbf{h}_{0}$. If we want to generate the corresponding parameters, we can then just use the second part of previous equation, namely,
@@
    Independent modes: Finally, the use of classifier-free guidance (CFG) [69] requires the definition of unconditional marginals, i.e. $p(\mathbf{h}_{0})$ and $p(\mathbf{w}_{0})$, which here we can sampled using
@@
As (ii) requires double the number of model evaluations, we only use (i) for joint sampling in Section˜IV.
@@ A.2 Multimodal top-down posterior
We show here that the top-down posterior from Eq.˜12 splits to Eq.˜13. Considering $0\leq s<t\leq 1$ and $0\leq\tilde{s}<\tilde{t}\leq 1$, we analyze the multimodal posterior by splitting the joint distribution
@@
In the previous equation, the second equality was done by means of the forward process Eq.˜3, that allows us to derive
@@
A similar derivation can be done for the term $q({\color[rgb]{0,1,1}\mathbf{h}_{t}}|\mathbf{h}_{s},\mathbf{w}_{\tilde{t}},\mathbf{x})$ in the denominator in Eq.˜20. We are using similar variants of Section˜A.2 below, i.e. for $q({\color[rgb]{0,.5,.5}\mathbf{h}_{s}}|{\color[rgb]{1,0,1}\mathbf{w}_{\tilde{s}}},\mathbf{w}_{\tilde{t}},\mathbf{x})$ and $q({\color[rgb]{0,.5,.5}\mathbf{h}_{s}}|\mathbf{w}_{\tilde{t}},\mathbf{x})$ in Eq.˜22, and for $q({\color[rgb]{0.76,0.01,0}\mathbf{w}_{\tilde{t}}}|\mathbf{h}_{t},{\color[rgb]{0,.5,.5}\mathbf{h}_{s}},\mathbf{x})$ and $q({\color[rgb]{0.76,0.01,0}\mathbf{w}_{\tilde{t}}}|\mathbf{h}_{t},\mathbf{x})$ in Eq.˜23 . Further, we simplify Eq.˜20 by
@@
which is the usual top-down posterior for $\mathbf{w}_{\tilde{s}}$ for standard Gaussian DMs, as used for instance in DDPM [32]. In the same direction as done above, for the top-down posterior of $\mathbf{h}_{s}$ (the first part of Eq.˜19) we derive
@@
The Eq.˜24 proves the splitting of the reverse transitions of Eq.˜13. Specifically, we set for the model
@@ A.3 Multimodal ELBO
which reduces to the loss of the main paper (see Eq.˜4) when accounting for: 1) all sampling modes from Section˜A.1; 2) the use of the multimodal top-down posterior of Section˜A.2.
@@ A.4 Velocity loss weight
As described in Section˜III.1, we use the velocity target [44]
@@
ensuring that the model outputs have unit variance. In contrast, the rectified flow-based velocity [70], defined as $\mathbf{v}_{\text{flow}}\coloneq{\bf\it\epsilon}-\mathbf{x}$, does not guarantee this property. For a Gaussian diffusion process (see Eq.˜2), we sample a latent $\mathbf{z}_{t}$ at time $t$ using
@@
where ${\bf\it\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Next, we combine Eq.˜28 and Eq.˜29 to
@@
Next, we derive the conversion factor $\omega_{x}(t)=(\mathrm{SNR}(t)+1)\,\omega_{v}(t)$, mentioned in Section˜III.1, that arises when one translates the loss from a denoising prediction $\hat{\mathbf{x}}_{\theta}$ to a velocity predictor $\hat{\mathbf{v}}_{\theta}$. For this, we insert Eq.˜30 into the MSE
@@
$\coloneq\omega_{v}(t)\norm{\left(\mathbf{v}_{t}-\hat{\mathbf{v}}_{\theta}\right)}_{2}^{2},$ (31)
@@
which we use for the loss weights in Eq.˜4, as introduced in the main text.
@@ A.5 Signal-to-noise ratio for continuous parameters
where $\bar{\beta}_{t}^{w}=1-\bar{\alpha}_{t}^{w}$. Moreover, we consider an embedding of the parameter $\lambda\in[-1,1]$ via (see Section˜III.2 and Section˜B.2)
@@
As explained in Section˜III.1, we use a sigmoid log-SNR weighting from Ref. [43]
@@
Notably, due to the encoding of the parameter $\lambda$ (see Eq.˜34 and Section˜B.2), we have to alter the normal SNR in Eq.˜35, initially given by $\mathrm{SNR}_{w}(t)\coloneq\bar{\alpha}_{t}^{w}/(1-\bar{\alpha}_{t}^{w})$. This redefinition is required because the signal is now encoded into two correlated dimensions, i.e. $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, and is therefore more robust to the noise of Eq.˜33. The new $\mathrm{SNR}_{\lambda}$ for the continuous parameters $\lambda$, with its noisy reconstruction $\hat{\lambda}_{t}$ (see Section˜B.2), is defined as
@@
Here, we assume that the samples of $\lambda$ from the prior $p(\lambda)$ are bounded and normalized. Since $\lambda$ is periodic in [1,1][-1,1] the absolute value is not relevant and we treat $\operatorname{\mathbb{E}}\left[\lambda^{2}\right]$ as a constant, as it does not depend on the diffusion time $t$. Indeed, we can view it as a constant shift for the sigmoid weighting in Eq.˜35.
Given the previous definition, we can get an upper bound of Eq.˜36 by using the Cramér–Rao bound [71] for an estimator of the parameter $\lambda$,
@@
where $\mathcal{I}(\cdot)$ is the Fisher information. We do so by first, calculating the score function $s(\lambda,\mathbf{w}_{t})$ of Eq.˜33
@@
with evaluating the derivative of Eq.˜34
@@
[…]
@@
where we used our orthogonal unit-variance normalized basis for $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ (see Section˜B.2). Given the previous, we define
@@
Therefore, for velocity prediction (see Section˜A.4), we set the weighting to
$\omega_{w}(t)$ $\coloneq(1-\bar{\alpha}_{t}^{w})\cdot\mathrm{sigmoid}[\log(\mathrm{SNR}_{\lambda}(t))]$
@@ A.6 Learned noise schedule for discrete tokens
The problem of too-weak-noise schedulers for diffusing discrete token embeddings is visualized in Fig.˜2a, e.g. for the cosine schedule [39], typically used for images. As explained in Section˜III.2, we aim to learn a noise schedule with a desired average Hamming distance profile, such that the time of trivial denoising is minimized. Note that the Hamming distance for a single token can only be either 0 (not flipped) or 1 (flipped). Instead, when averaging w.r.t. to a sampling process, we can use the average Hamming distance, which is effectively the probability of a token flipping into any other token.
@@
[…]
where $\tau>0$ is a temperature, which we generally set to $\tau=1/\sqrt{d_{h}}$ for a token embedding dimension of $d_{h}$. Analogous to the average Hamming distance, this probability $p_{\text{flip}}(t)$ is upper bounded by 11/N1-1/N, i.e., when $\mathbf{h}_{t}^{(i)}$ is sampled from the uniform distribution (see $t=1$ at Fig.˜2a).
@@
In order to match $p_{\text{flip}}(t)$ to a desired Hamming distance target $f_{\text{target}}(t)$, we optimize the noise schedule $\bar{\alpha}_{t}^{h}$ appearing in Section˜A.6 by minimizing
@@
In practice, we found it easier to optimize $\bar{\alpha}_{t}^{h}$ indirectly by parameterizing it with $\alpha_{t}^{h}$, using the cumulative product relation $\bar{\alpha}_{t(i)}^{h}=\prod_{k=0}^{i}\alpha_{t(k)}^{h}$ (see Eq.˜9 and Eq.˜10) and treating $\alpha_{t(k)}^{h}$ as learnable parameters. Before optimization, we initialize the schedule values $\alpha_{t(k)}^{h}$ with the cosine schedule [39].
@@
In Fig.˜2a we show the average Hamming distance of different optimized schedules according to Eq.˜54. The curves plotted correspond to the following noise schedules:
@@
In Section˜IV, we use for all experiments a learned discrete noise schedule for the linear target (see Fig.˜2a and c).
@@
As shown in Section˜III.3, we use the duality between uniform discrete-state diffusion and continuous-state Gaussian diffusion [46] to write the probability of finding a decoded, one-hot encoded token $\mathbf{k}_{t}\in\mathbb{R}^{N}$ for a discrete schedule $a_{t}$ as
@@
Just as the continuous SNR (see Eq.˜2), we have the same boundary behavior of no noise at small diffusion times ($\lim_{t\rightarrow 0}\mathrm{SNR}_{\text{discrete}}(t)\rightarrow\infty$) to full noise at time $t=1$ ($\mathrm{SNR}_{\text{discrete}}(t=1)=0$). Here, full noise means a uniform distribution across the classes.
Similar to the continuous mode, we use the sigmoid log-SNR weighting of Eq.˜35 (see Section˜III.1). In order to account for the discrete nature of the embeddings, which embody high correlations across the embedding dimensions, we adjust the normal $\mathrm{SNR}_{h}(t)$ (Section˜III.1) to the discrete one $\mathrm{SNR}_{\text{discrete}}(t)$ of Eq.˜57.
To determine the appropriate weight $\omega_{h}(t)$ in Eq.˜4, we relate the Gaussian flip probabilities $p_{\text{flip}}(t)$ (Section˜A.6) to the discrete schedule $a_{t}$ of Eq.˜56. As all tokens are equivalently encoded, it suffices to look at only one token $i$. Hence, following Eq.˜56, we write the probability of finding the token not to be flipped as
@@
resulting in the claimed relation of Section˜III.3
@@
Therefore, assuming a well optimized learned schedule with Hamming target $f_{\text{target}}(t)$ (see Eq.˜54), we can rewrite the discrete SNR as
@@
$\approx\frac{f_{\text{target}}(1)-f_{\text{target}}(t)}{f_{\text{target}}(t)}.$ (60)
Finally, for the velocity prediction (see Section˜A.4), we set the weighting to
$\omega_{h}(t)$ $\coloneq(1-\bar{\alpha}_{t}^{h})\cdot\mathrm{sigmoid}[\log(\mathrm{SNR}_{\text{discrete}}(t))]$